CURRICULUM VITAE

Endre Csáki

Date of Birth:  1935 January 7
Education, Degrees:  1958: M.Sc. in Mathematics

Eötvös Loránd University, Budapest

1975: Ph.D. (CSc) in Mathematics

Hungarian Academy of Sciences, Budapest

1989: DSc in Mathematics

Hungarian Academy of Sciences, Budapest

Positions:  1958-1975: Research Fellow, Mathematical Institute of the

Hungarian Academy of Sciences.

1970-1973: Lecturer, Department of Mathematics

Nairobi University (on leave from Math. Inst. Hung. Acad. Sci.)

1975-1989: Senior Research Fellow, Mathematical Institute

of the Hungarian Academy of Sciences.

1989-  : Scientific Advisor, Alfréd Rényi Institute of

Mathematics, Hungarian Academy of Sciences.

1985-2004: Head of Department of Probability at the

Mathematical Institute

A selection of important  1978 July: McMaster Univ, Canada
scientific visits  1981 May-June: Purdue Univ., USA
1981 July-August: McMaster Univ., Canada
1984 July-August: McMaster Univ., Canada
1990 July-August: Carleton Univ., Canada
1992 July-August: Carleton Univ., Canada
1997 August: Carleton Univ., Canada
1998 September: University Paris VI, France
1999 March and November: University Paris VI, France
Honours:  1964: Grünwald Géza Prize

awarded by Bolyai Math. Society

2003: Academic Prize

awarded by Hungarian Academy of Sciences

Memberships:  János Bolyai Mathematical Society, Hungary

American Mathematical Society Bernoulli Society

Personal:  married, father of three children